OpenStudy (anonymous):

How can x + 2 = -3x + 11 be set up as a system of equations?

OpenStudy (anonymous):

y=x+2 y=-3x+11 There's your system.

OpenStudy (anonymous):

My choices are : y = x2 - 9 y = x2 + 2x + 9 y = x2 y = x2 + 2x + 9 y = x2 + 2x y = x2 + 9 y = x2 y = 2x + 9

OpenStudy (anonymous):

@grazes

OpenStudy (anonymous):

In that case, I'm not sure what to make of it.